# Using Deep Learning to Design High Aspect Ratio Fusion Devices

### v1

@@ Abstract
The design of fusion devices is typically based on computationally expensive simulations. This can be alleviated using high aspect ratio models that employ a reduced number of free parameters, especially in the case of stellarator optimization where non-axisymmetric magnetic fields with a large parameter space are optimized to satisfy certain performance criteria. However, optimization is still required to find configurations with properties such as low elongation, high rotational transform, finite plasma beta, and good fast particle confinement. In this work, we train a machine learning model to construct configurations with favorable confinement properties by finding a solution to the inverse design problem, that is, obtaining a set of model input parameters for given desired properties. Since the solution of the inverse problem is non-unique, a probabilistic approach, based on mixture density networks, is used. It is shown that optimized configurations can be generated reliably using this method.
@@ 1 Introduction
Since analytically inverting the equations in the near-axis method is not feasible due to their differential integral character, a practical solution involves employing a machine learning model, specifically a neural network as a universal approximator (Hornik et al., 1989) to tackle the inverse problem. By training on a dataset of near-axis configurations, the neural network can learn either a forward mapping, from design parameters to configuration properties, or an inverse mapping, from configuration properties to design parameters. However, this inverse problem is ill-posed as multiple sets of design parameters can yield the same configuration properties, i.e., the standard stellarator design formulation is not bijective. Similarly to inverse design problems, using a neural network could lead to predicting an average of the design parameters for given configuration properties, resulting in inaccurate outcomes (Bishop, 1994). To overcome this challenge, we approximate the probability distribution of the design parameters conditioned on the configuration properties. This distribution, which can be multimodal (McLachlan & Basford, 1988), allows us to sample design parameters based on the desired configuration properties. To achieve this, a probabilistic machine learning model, namely the Mixture Density Networks (MDNs) model (Bishop, 1994), is used to solve the inverse problem of stellarator optimization, together with the near-axis expansion method. [^3] The code developed during this work is available at https://github.com/pedrocurvo/MLStellaratorDesign
@@ 2 Physical Model
The method for generating stellarator configurations in a near-axis expansion complements traditional stellarator optimization, which typically involves parameterizing the boundary shape of a finite aspect ratio plasma and using a 3D MHD equilibrium code to evaluate the objective function. Instead, in the near-axis method, we parameterize the axis curve and find the Taylor series coefficients of B in powers of $r$ that allow for quasisymmetry (Garren & Boozer, 1991b; Landreman & Sengupta, 2019; Jorge et al., 2020). While the near-axis method is necessarily approximate, it is orders of magnitude faster than standard methods, with a reduced parameter space, therefore allowing for broader parameter scans. Ultimately combining both approaches can be advantageous: the near-axis method can identify viable configurations, which can then be refined through conventional optimization.
@@
$\beta$ Plasma beta $2\mu_{0}p/B^{2}$.
@@
Although the near-axis expansion substantially reduces the number of free parameters, the optimization process may be computationally expensive, depending on the target parameters. Furthermore, it is necessary to compute the mapping between Boozer and Cartesian coordinates for a given surface to identify if the configuration possesses self-intersecting surfaces, making the process both time-consuming and resource-intensive. For this work, a viable configuration is one that meets the specific criteria outlined in Table 3. Such parameters are similar to the ones outlined in Landreman (2022). Informally, we will refer to stellarators that meet these criteria as good stellarators, and those that do not as bad stellarators.
@@ 4 Mixture Density Networks
By choosing a mixture model with a large enough number of components, and a neural network with a large enough number of hidden units, the MDN can approximate any conditional density $p(y|x)$ as closely as desired. In this work, we use a mixture model with 62 components. This choice was empirically determined to provide an optimal balance between model complexity and performance. It was observed that increasing both the number of layers and the width of each layer, as well as incorporating more components, provided severe improvements in the model’s performance. The architecture of the mixture density network used in this work is illustrated in Fig. 3 and in Table 4, showcasing the detailed configuration and activation functions employed at various layers.
@@ 5 Data Generation and Training
To generate the dataset, we sample the input parameters from uniform distributions, with the ranges listed in Table 5 and find the output parameters listed in Table 2. By sampling the input parameters from uniform distributions, we find that most configurations consist of bad stellarators. In fact, by applying the set of criteria shown in Table 3, it is seen that the percentage of good stellarators is extremely low, with only 1 in approximately 100,000 samples found to comply with all the desired criteria. This illustrates how difficult it is to find good stellarators by random search, and is one of the main drivers of the use of an inverse model to find the input parameters from a set of desired properties.
@@
Despite using the Adam optimizer (Kingma & Ba, 2017), the training process was more challenging than anticipated due to numerical instabilities, such as vanishing gradients, that caused the covariance matrices to become non-positive definite. To address this issue, a learning rate schedule was employed. The network was initially trained with a higher learning rate and then fine-tuned with a lower learning rate. This learning rate was decreased multiple times using a scheduler during the training until the network converged to the data distribution. The loss and validation curves can be seen in Fig. 4. Notably, the curves indicate that as the learning rate decreases, the loss function values also decrease. This trend suggests that lower learning rates contribute to a more stable and gradual convergence, resulting in better model performance and lower loss.
@@ 6 Model Performance
We now show how the model can be used to predict the input parameters needed to obtain optimized stellarators with desired output properties. First, the user provides the desired properties such as plasma $\beta$ and rotational transform, and the model produces the design parameters that are likely to yield those properties such as magnetic axis and $\bar{\eta}$. Then, the user feeds the predicted design parameters to the near-axis expansion method, to generate the corresponding properties. Finally, the user verifies that the actual properties generated by the near-expansion method agree with the desired properties. An example is shown in Table 7.
@@
Next, keeping the standard normalization on the dataset, we employed the Huber Loss and the Mean Absolute Error (MAE) as evaluation metrics to compare the predicted output properties from the model against the output properties from the near-axis model on 10,000 samples. Both are metrics used in regression tasks to quantify the difference between predicted values and actual observations. Huber Loss combines the advantages of MAE for robustness to outliers and Mean Squared Error (MSE) for sensitivity to small errors. The results for viable stellarators are presented in Table 8.
Variable Metric
Huber Loss MSE
axis length 0.031 0.0618
$\iota$ 0.0267 0.0534
max elongation 0.000456 0.0113
min $\text{min }L_{\nabla B}$ 0.266 0.735
min $\text{min }R_{0}$ 0.00617 0.0531
$r_{\text{singularity}}$ 0.632 1.72
$L_{\nabla\nabla B}$ 0.432 1.12
$B_{20_{\text{variation}}}$ 0.000604 0.0046
$\beta$ 0.321 1.1
$D_{\text{Merc}}\times r^{2}$ 3.4 ×$\times 10^{-11}$ 5.94 ×$\times 10^{-5}$
Average 0.172 0.486
Table 8: Model accuracy on viable stellarators
As illustrated in Table 8, the model accuracy was found to be satisfactory. For the variables axis length, $\iota$, max elongation, $B_{20_{\text{variation}}}$, and $D_{Merc}\times r^{2}$, the model showed a good performance, evidenced by a low Huber and MSE losses, 0.172 and 0.486 respectively, with the MSE being higher than the Huber Loss, as expected. Regarding the variables min $\text{min }L_{\nabla B}$, min $\text{min }R_{0}$, and $L_{\nabla\nabla B}$, the model displayed moderate accuracy under the Huber Loss metric. However, the MSE was higher, indicating that the model underperforms in these variables. The variables $\beta$ and $r_{\text{singularity}}$ exhibited the poorest accuracy, with both metrics indicating suboptimal results.
@@
For the viable stellarators, the distribution of the number of field periods, $n_{\text{fp}}$, is more narrowly centered around the value of 3, and none of the configurations exhibit $n_{\text{fp}}$ values above 6. This suggests a more constrained and specific range for $n_{\text{fp}}$ in the viable stellarator subset, indicating that these configurations are more consistent in this regard.
@@ 7 Conclusions
Despite achieving good performance in predicting some variables, the model faced challenges with variables derived from the second-order near-axis expansion method, as assessed using Huber Loss and Mean Absolute Error (MAE) metrics. Nevertheless, overall, the MDN proved effective as a tool for predicting desired properties of stellarators.
@@
As a future work, we intend to integrate the near-axis expansion method directly into the neural network training process, potentially as a differentiable layer. This advancement could leverage techniques like neural network approximations or automatic differentiation tools such as JAX (Bradbury et al., 2018). Such enhancements would support the adoption of variational autoencoders, graph neural networks, and transformers. Such models could also be extended for future optimizations and designs, integrating them with an ideal MHD model rather than relying solely on a near-axis method. Additionally, a model could be developed to map between the near-axis method and an ideal MHD model. This approach would enable leveraging machine learning models for solving the inverse problem using a near-axis method and subsequently mapping the results to a full ideal MHD optimization.
@@ Acknowledgments
We would like to thank Raheem Hashmani and Misha Padidar for their insightful discussions throughout this work. R. Jorge would like to acknowledge the support of EUROfusion through an Enabling Research Grant, and the support of FCT – Fundação para a Ciência e Tecnologia, I.P. through project reference 2021.02213.CEECIND/CP1651/CT0004. This work used Jetstream2 at Indiana University through allocation PHY240054 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, which is supported by National Science Foundation grants #213859, #2138286, #2138307, #2137603 and #2138296. This research used resources of the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231 using NERSC award NERSC DDR-ERCAP0030134. This research used resources of the Oak Ridge Leadership Computing Facility at the Oak Ridge National Laboratory, which is supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC05-00OR22725.

### v2

@@ Abstract
The design of fusion devices is typically based on computationally expensive simulations. This can be alleviated using high aspect ratio models that employ a reduced number of free parameters, especially in the case of stellarator optimization where non-axisymmetric magnetic fields with a large parameter space are optimized to satisfy certain performance criteria. However, optimization is still required to find configurations with properties such as low elongation, high rotational transform, finite beta, and good fast particle confinement. In this work, we train a machine learning model to construct configurations with favorable confinement properties by finding a solution to the inverse design problem, that is, obtaining a set of model input parameters for given desired properties. Since the solution of the inverse problem is non-unique, a probabilistic approach, based on mixture density networks, is used. It is shown that optimized configurations can be generated reliably using this method.
@@ 1 Introduction
Since analytically inverting the equations in the near-axis method is not feasible due to their differential integral character, a practical solution involves employing a machine learning model, specifically a neural network as a universal approximator (Hornik et al., 1989) to tackle the inverse problem. By training on a dataset of near-axis configurations, the neural network can learn either a forward mapping, from design parameters to configuration properties, or an inverse mapping, from configuration properties to design parameters. However, this inverse problem is ill-posed as multiple sets of design parameters can yield the same configuration properties (this was also observed in the database used in this work). This means that the standard stellarator design formulation is not bijective, in that it lacks a unique, one-to-one correspondence between design parameters and configuration properties. As with other inverse design problems, using a neural network in this context can result in predictions that represent an average of multiple possible design parameter sets for given configuration properties, rather than a specific solution. This averaging effect, described by Bishop (1994), can lead to inaccurate outcomes, as the network generalizes over multiple valid solutions instead of a unique parameter set. To overcome this challenge, we approximate the probability distribution of the design parameters conditioned on the configuration properties. This distribution, which can be multimodal (McLachlan & Basford, 1988), allows us to sample design parameters based on the desired configuration properties. To achieve this, a probabilistic machine learning model, namely the Mixture Density Networks (MDNs) model (Bishop, 1994), is used to solve the inverse problem of stellarator optimization, together with the near-axis expansion method. [^3] The code developed during this work is available at https://github.com/pedrocurvo/MLStellaratorDesign
@@ 2 Physical Model
The method for generating stellarator configurations in a near-axis expansion complements traditional stellarator optimization, which typically involves parameterizing the boundary shape of a finite aspect ratio plasma and using a 3D MHD equilibrium code to evaluate the objective function. The magnetic field equilibrium and plasma pressure are related via the ideal MHD equation $\mathbf{J}\times\mathbf{B}=\nabla p$ with $\mathbf{J}=\nabla\times\mathbf{B}/\mu_{0}$ the plasma current. Instead, in the near-axis method, we parameterize the axis curve and find the Taylor series coefficients of B in powers of $r$ that allow for quasisymmetry (Garren & Boozer, 1991b; Landreman & Sengupta, 2019; Jorge et al., 2020). While the near-axis method is necessarily approximate, it is orders of magnitude faster than standard methods, with a reduced parameter space, therefore allowing for broader parameter scans. Ultimately combining both approaches can be advantageous: the near-axis method can identify viable configurations, which can then be refined through conventional optimization.
@@
$\beta$ Volume-averaged plasma beta $\beta\right>=-\mu_{0}p_{2}r_{\text{singularity}}^{2}/B_{0}^{2}$.
@@
Although the near-axis expansion substantially reduces the number of free parameters, the optimization process may be computationally expensive, depending on the target parameters. Furthermore, it is necessary to compute the mapping between Boozer and Cartesian coordinates for a given surface to identify if the configuration possesses self-intersecting surfaces, making the process both time-consuming and resource-intensive. For this work, a viable configuration is one that meets the specific criteria outlined in Table 3. Such parameters are similar to the ones outlined in Landreman (2022). Those parameters also benefit the overall stability of the stellarators, e.g., $L_{\nabla B}$ is positively correlated with the coil-to-plasma distance, as demonstrated by Kappel et al. (2024), hence by constraining to larger values, we can obtain solutions with improved stability. Informally, we will refer to stellarators that meet these criteria as good stellarators, and those that do not as bad stellarators.
@@ 4 Mixture Density Networks
By choosing a mixture model with a large enough number of components, and a neural network with a large enough number of hidden units (Uzair & Jamil, 2020), the MDN can approximate any conditional density $p(y|x)$ as closely as desired  (Lu & Lu, 2020). In this work, we use a mixture model with 62 components. This choice was empirically determined to provide an optimal balance between model complexity and performance. It was observed that increasing both the number of layers and the width of each layer, as well as incorporating more components, provided severe improvements in the model’s performance. The architecture of the mixture density network used in this work is illustrated in Fig. 3 and in Table 4, showcasing the detailed configuration and activation functions employed at various layers.
@@ 5 Data Generation and Training
To generate the dataset, we sample the input parameters from uniform distributions, with the ranges listed in Table 5 and find the output parameters listed in Table 2. The range of parameters $R_{c2},R_{c3},Z_{c2}$ and $Z_{c3}$ follows the empirical observation in previous near-axis configurations and in the parameter scans done here that the Fourier coefficients generally decrease with increasing order. This allows us to restrict the database to feasible designs. By sampling the input parameters from uniform distributions, we find that most configurations consist of bad stellarators. In fact, by applying the set of criteria shown in Table 3, it is seen that the percentage of good stellarators is extremely low, with only 1 in approximately 100,000 samples found to comply with all the desired criteria. This illustrates how difficult it is to find good stellarators by random search, and is one of the main drivers of the use of an inverse model to find the input parameters from a set of desired properties.
@@
Despite using the Adam optimizer (Kingma & Ba, 2017), the training process was more challenging than anticipated due to numerical instabilities, such as vanishing gradients, that caused the covariance matrices to become non-positive definite. To address this issue, a multi-step learning rate scheduler was employed, which adjusted the learning rate at specific training epochs (10, 20, 30, 40, and 50) by a factor of 0.5. This schedule initially allowed the model to explore the parameter space with a higher learning rate, then gradually refined as training progressed. By reducing the learning rate in steps, the model avoided abrupt changes in parameter updates, leading to a more stable convergence. The loss and validation curves can be seen in Fig. 4. Notably, the curves indicate that as the learning rate decreases, the loss function values also decrease. This trend suggests that lower learning rates contribute to a more stable and gradual convergence, resulting in better model performance and lower loss.
@@ 6 Model Performance
We now show how the model can be used to predict the input parameters needed to obtain optimized stellarators with desired output properties. First, the user provides the desired properties such as the volume-averaged plasma $\beta$ and rotational transform, and the model produces the design parameters that are likely to yield those properties such as magnetic axis and $\bar{\eta}$. Then, the user feeds the predicted design parameters to the near-axis expansion method, to generate the corresponding properties. Finally, the user verifies that the actual properties generated by the near-expansion method agree with the desired properties. A randomly selected example from the dataset is presented in Table 7, while an example using the frontier conditions from Table 3 is shown in Table 8.
@@
Desired properties Design parameters Actual properties
(input to MDN) (output of MDN / (output of pyQSC)
input to pyQSC)
axis length 0.00 $R_{c1}$ 0.614602 7.99
$\iota$ 0.20 $R_{c2}$0.058358 -0.79
max elongation 10.0 $R_{c3}$ 0.020746 14.65
min $\text{min }L_{\nabla B}$ 0.10 $Z_{s1}$ 0.804627 0.21
min R0 0.30 $Z_{s2}$ 0.013770 0.31
$R_{\text{singularity}}$ 0.05 $Z_{s3}$ 0.013673 0.03
$L_{\nabla\nabla B}$ 0.10 $n_{\text{fp}}$ 1 0.10
$B_{20_{\text{variation}}}$ 5.00 $\bar{\eta}$ 0.509180 5.51
$\beta$ 0.001 B2c -1.257980 0.00014
$D_{\text{Merc}}\times r^{2}$ 0 p2 -154446 0.10
Table 8: Sample results for given desired properties that were the boundary conditions in Table 3. The properties of the given stellarator were used as input to the model to predict the design parameters. These predicted design parameters were then fed into the Near-Axis Method, which returned the actual properties. The resulting actual properties closely matched the desired ones.
@@
Next, keeping the standard normalization on the dataset, we employed the Huber Loss and the Mean Absolute Error (MAE) as evaluation metrics to compare the predicted output properties from the model against the output properties from the near-axis model on 10,000 samples. Both are metrics used in regression tasks to quantify the difference between predicted values and actual observations. Huber Loss combines the advantages of MAE for robustness to outliers and Mean Squared Error (MSE) for sensitivity to small errors. The results for bad, good and viable stellarators are presented in Table 9.
Variable Viable Good Bad
Metric Metric Metric
Huber Loss MSE Huber Loss MSE Huber Loss MSE
axis length 0.031 0.0618 0.0342 0.083 1.33 10.2
$\iota$ 0.0267 0.0534 0.0233 0.0495 0.909 4.2
max elongation 0.000456 0.0113 0.000326 0.068 0.138 17.8
min $\text{min }L_{\nabla B}$ 0.266 0.735 0.227 0.578 0.869 3.82
min $\text{min }R_{0}$ 0.00617 0.0531 0.00665 0.274 2.72 32.6
$r_{\text{singularity}}$ 0.632 1.72 0.907 3.11 0.0098 0.0292
$L_{\nabla\nabla B}$ 0.432 1.12 0.415 1.08 0.149 0.366
$B_{20_{\text{variation}}}$ 0.000604 0.0046 8.99×$\times 10^{-5}$ 0.000218 3.9 37.4
$\beta$ 0.321 1.1 0.658 3.12 0.00454 0.0165
$D_{\text{Merc}}\times r^{2}$ 3.4 ×$\times 10^{-11}$ 5.94 ×$\times 10^{-5}$ 1.65×$\times 10^{-10}$ 9.94×$\times 10^{-12}$ 83.7 124.3
Average 0.172 0.486 0.227 0.837 9.37 23.073
Table 9: Model accuracy on bad, good and viable stellarators
As illustrated in Table 9 for viable stellarators, the model accuracy was found to be satisfactory. For the variables axis length, $\iota$, max elongation, $B_{20_{\text{variation}}}$, and $D_{Merc}\times r^{2}$, the model showed a good performance, evidenced by a low Huber and MSE losses, 0.172 and 0.486 respectively, with the MSE being higher than the Huber Loss, as expected. Regarding the variables min $\text{min }L_{\nabla B}$, min $\text{min }R_{0}$, and $L_{\nabla\nabla B}$, the model displayed moderate accuracy under the Huber Loss metric. However, the MSE was higher, indicating that the model underperforms in these variables. The variables $\beta$ and $r_{\text{singularity}}$ exhibited the poorest accuracy, with both metrics indicating suboptimal results. A possible explanation for this outcome might be due to trade-offs in variable correlations, i.e., maximizing performance for some variables may require sacrificing accuracy in others.
@@
For the viable stellarators, the distribution of the number of field periods, $n_{\text{fp}}$, is more narrowly centered around the value of 3, and none of the configurations exhibit $n_{\text{fp}}$ values above 6. This suggests a more constrained and specific range for $n_{\text{fp}}$ in the viable stellarator subset, indicating that these configurations are more consistent in this regard. The fact that an optimized stellarator with a higher number of field periods is hard to find, as it was also observed in Landreman (2022), may be related to the fact that such $n_{\text{fp}}$ usually require a significant excursion of the axis and an associated larger axis length. Furthermore, the recent study by Kappel et al. (2024) has shown a correlation between the number of field periods $n_{p}$ and $L_{\nabla B}$, indicating that small values of $n_{p}$ may lead to more optimized configurations.
@@ 7 Conclusions
Despite achieving good performance in predicting some variables, the model faced challenges with variables derived from the second-order near-axis expansion method, as assessed using Huber Loss and Mean Absolute Error (MAE) metrics. Nevertheless, overall, the MDN proved effective as a tool for predicting desired properties of stellarators. Our model can also return the covariance matrix to compute uncertainties associated with each prediction and obtain statistical insight.
@@
As a future work, an ablation study would be crucial to simplify the model, as the increasing complexity of the hidden layer geometry may not be optimal. Adding to this, we intend to integrate the near-axis expansion method directly into the neural network training process, potentially as a differentiable layer. This advancement could leverage techniques like neural network approximations or automatic differentiation tools such as JAX (Bradbury et al., 2018). Such enhancements would support the adoption of variational autoencoders, graph neural networks, and transformers. Such models could also be extended for future optimizations and designs, integrating them with an ideal MHD model rather than relying solely on a near-axis method. Additionally, a model could be developed to map between the near-axis method and an ideal MHD model. This approach would enable leveraging machine learning models for solving the inverse problem using a near-axis method and subsequently mapping the results to a full ideal MHD optimization.
@@ Acknowledgments
We would like to thank Raheem Hashmani and Misha Padidar for their insightful discussions throughout this work. R. Jorge would like to acknowledge the support of EUROfusion through an Enabling Research Grant, and the support of FCT – Fundação para a Ciência e Tecnologia, I.P. through project reference 2021.02213.CEECIND/CP1651/CT0004. This work has been carried out within the framework of the EUROfusion Consortium, funded by the European Union via the Euratom Research and Training Programme (Grant Agreement No 101052200 — EUROfusion). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Commission. Neither the European Union nor the European Commission can be held responsible for them. This work used Jetstream2 at Indiana University through allocation PHY240054 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, which is supported by National Science Foundation grants #213859, #2138286, #2138307, #2137603 and #2138296. This research used resources of the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231 using NERSC award NERSC DDR-ERCAP0030134. This research used resources of the Oak Ridge Leadership Computing Facility at the Oak Ridge National Laboratory, which is supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC05-00OR22725.